OpenStudy (anonymous):

SOLVE THE EQUATION BY COMPLETING THE SQUARE. X SQUARE +8x-33=0

OpenStudy (anonymous):

\[x ^{2}+8x-33\]

OpenStudy (anonymous):

First, take the number infront of x, divide it by 2, and then square it.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Alright, now add and then subtract it after 8x so you have: \[x ^{2}+8x+16-16-33\]

OpenStudy (anonymous):

and then factor the \[x ^{2}+8x+16\]

OpenStudy (anonymous):

Hint: itll be (x-a)^2 where a equals the number you divided by 2 before you squared it

OpenStudy (anonymous):

so 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So now we have (x+4)^2-16-33

OpenStudy (anonymous):

Move the numbers to the right and you get (x+4)^2=49

OpenStudy (anonymous):

What would you do next to solve for x?

OpenStudy (anonymous):

ok -49 add

OpenStudy (anonymous):

wait -4

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

If you're solving for x you would take the square root of both sides. it cancels the square out on the left so we have \[x+4=\pm \sqrt{49}\]

OpenStudy (anonymous):

then it would be7, -1?

OpenStudy (anonymous):

square root 7

OpenStudy (anonymous):

No, now you subtract the 4 from both sides and end up with: \[x=-4\pm7\]

OpenStudy (anonymous):

Since 7 is the square root of 49.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Since its plus or minus, you have two separate answers. x=-4+7 and x=-4-7 or x=3 and x=-11

OpenStudy (anonymous):

But when you take a square root, you have to go back and check your answers. So plug 3 in for x in the original equation and if its true, that is an answer. Same for -11

OpenStudy (anonymous):

ok that would be an easier way of checking.

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

Did you check the answers? in this case both 3 and -11 are solutions. Post your question as a new one and i will look at it

OpenStudy (anonymous):

ok